# Spectra of Magnetoroton and Chiral Graviton Modes of Fractional Chern Insulator

### v1

@@ Abstract
FCIs have been studies in many aspects in connection with FQHE, including the similarity and difference between topological flat bands and Landau levels [20, 21, 22, 23], the correspondence between their wave functions [24] and the ground state entanglement and generalized Pauli principle [6]. As for the collective excitations, the charge-neutral excitations are known to play a dominate role in controlling the properties of the FQHE/FCI systems. For example, the softening of magnetoroton (minimum) is an intrinsic mechanism to trigger phase transitions from FQHE/FCI to CDW [25, 26, 27, 28, 29, 30, 4, 31, 32, 33, 34]. The magnetoroton in the long-wavelength limit is known as the chiral graviton [35], reflecting the oscillation of the geometrical properties, which carries the spin angular momentum $2$ [36, 37, 38, 39, 40, 41]. This geometrical (apart from the topological) description of the FQHE was proposed by Haldane [42] and the experimental observation of chiral gravitons of FQH states was recently reported [43]. Similar as the roton minimum, the graviton mode could also approach the ground-state energy and leads to the FQH nematic state with anisotropic transport [44, 45, 46, 47].
@@
Before discussing the spectra, we begin with the ground state phase diagram as shown in Fig.1, The system go through a FCI-Solid I-Solid II phase transition triggered by the condensation of two successive rotons (with momenta $\mathbf{b_{1}}/2$ and $\mathbf{b_{1}}/4+\mathbf{b_{2}}/2$ respectively) as we increase $V_{2}$ along the path $V_{1}=4$. Such successive phase transition can be interpreted as a vestigial transition [58, 59, 60, 61] with the translational symmetry of the lattice broken in a two-step manner [29]. Here, static density structure factor $S^{A}(\mathbf{k})=\frac{1}{N_{A}}\sum_{ij}e^{-i\mathbf{k}\cdot(\mathbf{x_{i}-x_{j}})}\langle n_{i,A}n_{j,A}\rangle$ and charge order parameter $\delta^{A}(\mathbf{k})=\frac{1}{\sqrt{N_{A}}}\sum_{i}e^{-i\mathbf{k}\cdot\mathbf{x_{i}}}\langle n_{i,A}\rangle$ are measured to diagnosis the phases, where $A$ denotes the sub-lattice, $N_{A}$ denotes its site number, $i,j$ run over all unit cells. Since the results are the same for both sublattices, we mainly present the results on $A$ sublattice. The bright point at $\frac{\mathbf{b_{1}}}{2}$ in Fig. 1 (a) denotes the location of the lowest roton minimal of FCI [25, 26, 29], the fact that the rotational symmetry is not recovered in the $S^{A}(\mathbf{k})$ is a consequence of the cylindrical geometry of the DMRG simulation. In this system there is a second roton in the FCI phase, and it is located at the $\mathbf{b_{1}}/4+\mathbf{b_{2}}/2$ (it can be clearly visualised in the spectra in Fig. 2 (b)), and once the lowest roton condenses and the system enters the Solid I phase, the measurement of $S^{(A)}(\mathbf{k})$ in Fig. 1 (b) shows that the brightest density correlation is at this wave vector ($V_{1}=4,V_{2}=3.2$). The charge order of Solid I, however, can be seen from its order parameter $\delta^{A}(\mathbf{k})$ in Fig. 1 (e) ($V_{1}=4,V_{2}=3.2$). And the grey arrow connecting the momentum $\mathbf{b_{1}}/2$ from Fig. 1 (a) to Fig. 1 (e), symbolises the condensation of the roton and formation of the charge order. The process also happened in transition from Solid I to Solid II, symbolised by the grey arrow connecting the momentum $\mathbf{b_{1}}/4+\mathbf{b_{2}}/2$ of the second roton in Fig. 1 (b) to the charge order in Fig. 1 (f) ($V_{1}=4,V_{2}=4.6$).
@@
$S^{A}(\mathbf{k},\omega)$ $=\frac{1}{\sqrt{N_{t}N_{A}}}\sum_{jl}\ e^{i\omega t_{l}}e^{-i\mathbf{k}(\mathbf{x_{j}-x_{0}})}\big{(}\langle n_{j,A}(t_{l})n_{0,A}\rangle$ (2)
@@
where $n_{j,A}(t_{l})=e^{iHl\Delta t}n_{j,A}e^{-iHl\Delta t}$ as $l$ runs from 00 to $N_{t}$. We time evolve $n_{0,A}|\psi_{0}\rangle$ with $|\psi_{0}\rangle$ the ground state wave function and $n_{0,A}$ defined in the bulk of the cylinder [48]. The correlation function $\langle\hat{n}_{j,A}(t)\hat{n}_{0,A}\rangle$ is measured at every time slice as we progressively evolve $n_{0,A}|\psi_{0}\rangle$ with operator $e^{-iHt_{l}}$ in TDVP. For $N_{t}$, $n_{0,A}$ and the details of spectrum calculation, see SM [48].

### v2

@@ Abstract
FCIs have been studies in many aspects in connection with FQHE, including the similarity and difference between topological flat bands and Landau levels [20, 21, 22, 23], the correspondence between their wave functions [24] and the ground state entanglement and generalized Pauli principle [6]. As for the collective excitations, the charge-neutral excitations are known to play a dominate role in controlling the properties of the FQHE/FCI systems. For example, the softening of magnetoroton (minimum) is an intrinsic mechanism to trigger phase transitions from FQHE/FCI to CDW [25, 26, 27, 28, 29, 30, 4, 31, 32, 33, 34]. The magnetoroton in the long-wavelength limit is known as the chiral graviton [35], reflecting the oscillation of the geometrical properties, which carries the spin angular momentum $2$ [36, 37, 38, 39, 40, 41]. This geometrical (apart from the topological) description of the FQHE was proposed by Haldane [42] and the experimental observation of chiral gravitons of FQH states was recently reported [43]. Similar as the roton minimum, the graviton mode could also approach the ground-state energy which leads to the FQH nematic state with anisotropic transport [44, 45, 46, 47] and describe quench dynamics [48, 49].
@@
Before discussing the spectra, we begin with the ground state phase diagram as shown in Fig.1, The system go through a FCI-Solid I-Solid II phase transition triggered by the condensation of two successive rotons (with momenta $\mathbf{b_{1}}/2$ and $\mathbf{b_{1}}/4+\mathbf{b_{2}}/2$ respectively) as we increase $V_{2}$ along the path $V_{1}=4$. Such successive phase transition can be interpreted as a vestigial transition [60, 61, 62, 63, 64, 65] with the translational symmetry of the lattice broken in a two-step manner [29]. Here, static density structure factor $S^{A}(\mathbf{k})=\frac{1}{N_{A}}\sum_{ij}e^{-i\mathbf{k}\cdot(\mathbf{x_{i}-x_{j}})}\langle n_{i,A}n_{j,A}\rangle$ and charge order parameter $\delta^{A}(\mathbf{k})=\frac{1}{\sqrt{N_{A}}}\sum_{i}e^{-i\mathbf{k}\cdot\mathbf{x_{i}}}\langle n_{i,A}\rangle$ are measured to diagnosis the phases, where $A$ denotes the sub-lattice, $N_{A}$ denotes its site number, $i,j$ run over all unit cells. Since the results are the same for both sublattices, we mainly present the results on $A$ sublattice. The bright point at $\frac{\mathbf{b_{1}}}{2}$ in Fig. 1 (a) denotes the location of the lowest roton minimal of FCI [25, 26, 29], the fact that the rotational symmetry is not recovered in the $S^{A}(\mathbf{k})$ is a consequence of the cylindrical geometry of the DMRG simulation. In this system there is a second roton in the FCI phase, and it is located at the $\mathbf{b_{1}}/4+\mathbf{b_{2}}/2$ (it can be clearly visualised in the spectra in Fig. 2 (b)), and once the lowest roton condenses and the system enters the Solid I phase, the measurement of $S^{(A)}(\mathbf{k})$ in Fig. 1 (b) shows that the brightest density correlation is at this wave vector ($V_{1}=4,V_{2}=3.2$). The charge order of Solid I, however, can be seen from its order parameter $\delta^{A}(\mathbf{k})$ in Fig. 1 (e) ($V_{1}=4,V_{2}=3.2$). And the grey arrow connecting the momentum $\mathbf{b_{1}}/2$ from Fig. 1 (a) to Fig. 1 (e), symbolises the condensation of the roton and formation of the charge order. The process also happened in transition from Solid I to Solid II, symbolised by the grey arrow connecting the momentum $\mathbf{b_{1}}/4+\mathbf{b_{2}}/2$ of the second roton in Fig. 1 (b) to the charge order in Fig. 1 (f) ($V_{1}=4,V_{2}=4.6$).
@@
$S^{A}(\mathbf{k},\omega)$ $=\frac{1}{\sqrt{N_{t}N_{A}}}\sum_{jl}\ e^{i(\omega+i\eta)t_{l}}e^{-i\mathbf{k}(\mathbf{x_{j}-x_{0}})}\big{(}\langle n_{j,A}(t_{l})n_{0,A}\rangle$ (2)
@@
where $n_{j,A}(t_{l})=e^{iHl\Delta t}n_{j,A}e^{-iHl\Delta t}$ as $l$ runs from 00 to $N_{t}$ and $\eta=0.05$. We time evolve $n_{0,A}|\psi_{0}\rangle$ with $|\psi_{0}\rangle$ the ground state wave function and $n_{0,A}$ defined in the bulk of the cylinder [50]. The correlation function $\langle\hat{n}_{j,A}(t)\hat{n}_{0,A}\rangle$ is measured at every time slice as we progressively evolve $n_{0,A}|\psi_{0}\rangle$ with operator $e^{-iHt_{l}}$ in TDVP. For $N_{t}$, $n_{0,A}$ and the details of spectrum calculation, see SM [50].